Circular No. 9003

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 X3 (LINEAR) An apparently asteroidal object discovered by the LINEAR survey

(and posted on the Minor Planet Center’s ‘NEOCP’ webpage; discovery observation tabulated below) has been reported by CCD astrometrists elsewhere to show cometary appearance. R. A. Kowalski (Mt. Lemmon 1.5-m reflector) writes that four co-added 30-s CCD exposures taken on Dec. 5.5 UT in good seeing show a coma of diameter ≈ 15′ and a tail ≈ 30′′ long in p.a. 300^{o}. G. Hug (Scranton, KS, U.S.A., 0.56-m reflector) notes the object to be an “obvious comet” with a nebulous extension in p.a. ∼ 330^{o} on his images around the same time.

2008 UT  α_{2000}  δ_{2000}  Mag.
Dec. 4.44689  12^{h}20^{m}41.^{s}89  +8^{o}46′23′′.5  18.4

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-X48.

T = 2008 Nov. 19.300 TT  ω = 161.^{o}971
Ω = 338.258 }2000.0
q = 2.11036 AU  i = 71.042

COMETS C/2008 P2–P6 AND C/2008 Q5 (SOHO) Further to IAUC 9001, additional Kreutz sungrazers have been found

on SOHO website images. C/2008 P2 was small and stellar in appearance (mag ∼ 6.5). C/2008 P3 was small and stellar in appearance in C3 images (mag ∼ 6.5), and very diffuse, very faint, and elongated in C2 images. C/2008 P4 was small and stellar in appearance (mag ∼ 7). C/2008 P5 was small (mag ∼ 6.5) with a short tail in C3 images, and very diffuse, very faint, and elongated in C2 images. C/2008 P6 (found by Z. Xu = ZX) was stellar in appearance (mag ∼ 7). C/2008 Q5 was tiny, very faint (mag ∼ 7–7.5), and stellar in appearance.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 P2 Aug. 1.388 8^{h}25.^{m}9 +15^{o}40′ C3  HS 2008-T11
C/2008 P3 Aug. 6.196 8 43.2 +14 23 C3/2 MK 2008-T11
C/2008 P4 Aug. 8.054 8 53.4 +14 53 C3  MK 2008-T11
C/2008 P5 Aug. 9.654 8 57.7 +13 18 C3/2 TH 2008-T11
C/2008 P6 Aug. 14.446 9 19.8 +13 12 C3  ZX 2008-T11
C/2008 Q5 Aug. 26.763 10 06.2 + 9 10 C3  JR 2008-X24
2008 December 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green